# Localization landscape of optical waves in multifractal photonic membranes

### 05 Mar 2024 v2

@@ 3 The Helmholtz localization landscape of one-dimensional multifractals
Figure 1: Tight-binding optical potential of (a) single waveguide and (b) coupled waveguide system. (c,d) Spatial profiles of the lowest 3 optical modes from the solution of Eq. 5, along with the calculated landscape function $u$ (doted lines) computed from Eq. 7.
@@
In order to demonstrate the predictive power of the Helmholtz landscape introduced above, we consider first the simple case of a three-layer asymmetric waveguide structure with a dielectric permittivity in the core region $\epsilon_{c}=\epsilon_{max}=12.25$ and with cladding permittivities $\epsilon_{1}=1$ and $\epsilon_{2}=2.10$. We use a wavelength $\lambda=1.55\mu$m and display the tight-binding optical potential from Eq. 6 and Eq. 8 in Figure 1(a). In addition, we considered a system of two coupled dielectric waveguides with the optical potential shown in Figure 1(b). These are canonical eigenproblems in photonics whose solutions can easily be obtained with traditional methods [30]. Here, by computing the eigensolutions of the corresponding Schrödinger equation 5, we plot the first 3 optical modes of the two investigated systems in Figures 1(c,d), respectively. The dashed lines indicate the corresponding Helmholtz landscapes obtained by solving equation 7. In all the calculations we set the hopping rate to be $t=1$, without loss of generality. The results plotted in Figure 1 indicate that the Helmholtz landscape accurately predicts the localization regions of the optical modes computed from the tight-binding solution of the equivalent Schrödinger problem.
Figure 3: Typical eigenmodes of 1D multifractal structures generated by (a) the multiplicative cascade method and (b) the MMAR method. (c-d) Geometrical supports of the eigenmodes predicted by the landscape function and compared to the actual positions of the localized mode maxima, corresponding to the structures in (a) and (b).
@@ 4 The Helmholtz localization landscape of two-dimensional multifractals
Figure 5: (a-h) Eigenmodes whose spectral parameters vary in the range $0.80.55<\overline{d}_{1}/\lambda<$ 0.8. (i) Superposition of eigenmodes corresponding to a narrow range (within a $5\%$ dispersion value) of eigenvalues around $k_{0}^{2}$. Incident wavelengths $\lambda$ are indicated in each panel.
@@
Figure 6: (a-i) Eigenmodes whose spectral parameters vary in the range $0.780.6<\overline{d}_{1}/\lambda<$ 0.78. (i) Superposition of eigenmodes corresponding to a narrow range (within a $5\%$ dispersion value) of eigenvalues around $k_{0}^{2}$. Incident wavelengths $\lambda$ are indicated in each panel.
The landscape function $u(\mathbf{r})$ of both monofractal and multifractal structures are plotted in Figures 4(c,d). One can see that the landscape of the monofractal potential features intense maxima localized within small groups of locally-symmetric clusters of adjacent scattering particles distributed across the structure with spatial distributions that closely resemble its fractal support. These are the regions in which the localization landscape predicts the existence of highly confined resonant modes localized by proximity effects occurring among closely coupled particles. This is the typical behavior observed in small fractal aggregates of dipolar particles [37]. On the other hand, the landscape of the multifractal potential displays localization regions that are more broadly distributed across the support of the structure. This is consistent with the non-homogeneous multiscale nature of multifractal systems [20]. In Figures 5(a-h) and 6(a-h), we display the calculated eigenmodes at the wavelengths specified in each panel. As for the 1D simulations, particular care has been taken to ensure that the eigenvectors are narrowly distributed around $k_{0}^{2}$ (within a $5\%$ dispersion value). Moreover, in order to facilitate comparisons with the experimental measurements in which multiple modes are simultaneously excited by an external focused beam, we also plotted in Figures 5(i) and 6(i) a linear combination of the modes shown in panels (a-h). As predicted by the localization landscape, we obtain strongly localized modes within small particle clusters for the monofractal structure, which are typical of self-similar arrays where they have been successfully exploited to dramatically increase the cross sections of nonlinear and Raman processes [37]. However, the situation is quite different in the multifractal structure, where high-intensity localized modes appear to spread across a more extended region of the geometrical support, reflecting its larger degree of spatial non-uniformity compared to the ones of monofractals. This is evident by looking at the mode superposition in Figure 6(i) in which the amplitude peaks cover a larger fraction of the structure. In the next sections, we discuss the fabrication and characterization of dielectric membrane structures designed based on the localization landscape theory and we demonstrate experimentally the formation of characteristic fractal modes in qualitative agreement with the landscape predictions.
@@ 5.3 Experimental characterization of multifractal resonances
As shown in in Figure 10, the observed mode patterns of the multifractal structure display highly fluctuating spatial profiles distributed across the entire geometrical support, in qualitative agreement with the simulation results shown in Figure 6. Moreover, due to the spatially non-homogeneous nature of the investigated multifractal structure, the measured modes display no evident trend in the degree of their spatial localization when varying the wavelength, which is indicative of a broader spectrum of localized resonances compared to monofractals [20]. These results, demonstrate the distinctive localization behavior of the scattering resonances supported by large-scale fractal and multifractal photonic membranes designed by the efficient localization landscape approach. Future work will focus on the quantitative analysis of the spectral measures of the investigated multifractal optical systems, such as their integrated density of states [17].
@@ 6 Conclusion
In this paper, we proposed a method to efficiently design extended photonic structures with multifractal geometries and investigate the fundamental properties of the fractal eigenmodes of 1D and 2D Helmholtz operators with multifractal scattering potentials. In particular, we explored two canonical multifractal systems generated from the multiplicative cascade and the MMAR methods. Without solving the associated eigenproblems, we calculated the landscape functions and accurately predicted the locations of the supported eigenmodes. Finally, based on the information obtained from the localization landscapes, we designed and fabricated multifractal photonic membranes in highly transparent SiN materials and directly imaged, using leaky-mode spectroscopy, their optical modes across the visible spectral range. The general predictions from the localization landscape of the Helmholtz operator were found to be in good qualitative agreement with the experimental data, establishing the localization landscape theory as a viable tool for the rapid exploration and benchmarking of scattering resonances in complex photonic structures with tailored multifractal disorder for nanophotonics and metamaterials applications.

### 14 Mar 2024 v3

@@ 3 The Helmholtz localization landscape of one-dimensional multifractals
Figure 1: Tight-binding optical potential of (a) asymmetric slab waveguide and (b) coupled slab waveguides system. (c,d) Spatial profiles of the lowest three optical modes from the solution of Eq. 5 along with the calculated landscape function $u$ (doted lines) computed from Eq. 7.
@@
In order to demonstrate the predictive power of the Helmholtz landscape introduced above, we consider first the simple case of a three-layer asymmetric waveguide structure with a dielectric permittivity in the core region $\epsilon_{c}=\epsilon_{\mathrm{max}}=12.25$ and with cladding permittivities $\epsilon_{1}=1.00$ and $\epsilon_{2}=2.10$. We use a wavelength $\lambda=1.55\,\mu$m and display the tight-binding optical potential from Eq. 4 and Eq. 8 in Figure 1(a). In addition, we considered a system of two coupled dielectric waveguides with the optical potential shown in Figure 1(b). These are canonical eigenproblems in photonics whose solutions can easily be obtained with traditional methods [30]. Here, by computing the eigensolutions of the corresponding Schrödinger Eq. 5, we plot the first three optical modes of the two investigated systems in Figures 1(c,d), respectively. The dashed lines indicate the corresponding Helmholtz landscapes obtained by solving Eq. 7. In all the calculations we set the hopping rate to be $t=1$, without loss of generality. The results plotted in Figure 1 indicate that the Helmholtz landscape accurately predicts the localization regions of the optical modes computed from the tight-binding solution of the equivalent Schrödinger problem.
Figure 3: Correlation between the predicted and computed locations of the localized eigenmodes of 1D multifractal structures generated by (a) the multiplicative cascade method and (b) the MMAR method. The geometrical supports of the eigenmodes are predicted by the landscape function as explained in the text.
@@ 4 The Helmholtz localization landscape of two-dimensional multifractals
Figure 5: (a-h) Eigenmodes whose spectral parameters vary in the range $0.80.55<\overline{d}_{1}/\lambda<$ 0.8. (i) Superposition of eigenmodes corresponding to a narrow range (within a $5\%$ dispersion value) of eigenvalues around $k_{0}^{2}$. The wavelengths $\lambda$ at which the modes are computed are indicated in each panel.
@@
Figure 6: (a-i) Eigenmodes whose spectral parameters vary in the range $0.780.6<\overline{d}_{1}/\lambda<$ 0.78. (i) Superposition of eigenmodes corresponding to a narrow range (within a $5\%$ dispersion value) of eigenvalues around $k_{0}^{2}$. The wavelengths $\lambda$ at which the modes are computed are indicated in each panel.
The landscape function $u(\mathbf{r})$ of both monofractal and multifractal structures are plotted in Figures 4(c,d). One can see that the landscape of the monofractal potential features intense maxima localized within small groups of locally-symmetric clusters of adjacent scattering particles distributed across the structure with spatial distributions that closely resemble its fractal support. These are the regions in which the localization landscape predicts the existence of highly confined resonant modes localized by proximity effects occurring among closely coupled particles. This is the typical behavior observed in small fractal aggregates of dipolar particles [37]. On the other hand, the landscape of the multifractal potential displays localization regions that are more broadly distributed across the support of the structure. This is consistent with the non-homogeneous multiscale nature of multifractal systems [20]. In Figures 5(a-h) and 6(a-h), we display the calculated eigenmodes at the wavelengths specified in each panel. As for the 1D simulations, particular care has been taken to ensure that the eigenvectors are narrowly distributed around $k_{0}^{2}$ (within a $5\%$ dispersion value). Moreover, in order to facilitate comparisons with the experimental measurements in which multiple modes are simultaneously excited by an external source, we also plotted in Figures 5(i) and 6(i) a linear combination of the modes that independently contribute to the overall intensity distribution at $\lambda=720$ nm and $\lambda=750$ nm, respectively. As predicted by the localization landscape, we obtain strongly localized modes within small particle clusters for the monofractal structure, which are typical of self-similar arrays where they have been successfully exploited to dramatically increase the cross sections of nonlinear and Raman processes [37]. However, the situation is quite different in the multifractal structure, where high-intensity localized modes appear to spread across a more extended region of the geometrical support, reflecting its larger degree of spatial non-uniformity compared to the ones of monofractals. This is evident by looking at the mode superposition in Figure 6(i) in which the amplitude peaks cover a larger fraction of the structure. In the next sections, we discuss the fabrication and characterization of dielectric membrane structures designed based on the localization landscape theory and we demonstrate experimentally the formation of characteristic fractal modes in qualitative agreement with the landscape predictions.
@@ 5.3 Experimental characterization of multifractal resonances
As shown in in Figure 10, the observed mode patterns of the multifractal structure display highly fluctuating spatial profiles distributed across the entire geometrical support, in qualitative agreement with the simulation results shown in Figure 6. Moreover, due to the spatially non-homogeneous nature of the investigated multifractal structure, the measured modes display no evident trend in the degree of their spatial localization when varying the wavelength, which is indicative of a broader spectrum of localized resonances compared to monofractals [20]. We emphasize that, since the Helmholtz landscape method introduced here relies on scalar fields and point-like particles, it can only offer basic insights into the general nature of the localized modes. Specifically, these insights relate to the ability to pinpoint exactly the localization regions of the modes and predict global properties such as the localization/delocalization behavior of the resonant states. Therefore, it is unreasonable to expect a quantitative agreement with the experimental data at this stage. In particular, the model developed here does not capture the effects of the finite-size of the scattering particles, which may introduce local (small-scale) modifications to the overall mode structures compared to the ones measured on the fabricated samples. However, our work shows that the landscape approach is still valuable in providing predictive insights at low computational cost in large-scale scattering structures with complex potentials that are often beyond the reach of fully numerical, grid-based techniques such as the finite element method. Thus, we consider the proposed approach as a viable tool for the efficient first-order design of complex photonics structures. Future work will focus on the quantitative analysis of the spectral measures of the investigated multifractal optical systems, such as their integrated density of states [17], and the statistical distribution of intensity maxima of eigenvectors that unveil long-range multifractal correlations beyond the traditional Anderson model [45, 46, 47].
@@ 6 Conclusion
In this paper, we proposed a method to efficiently design extended photonic structures with multifractal geometries and investigate the fundamental properties of the fractal eigenmodes of 1D and 2D Helmholtz operators with multifractal scattering potentials. In particular, we explored two canonical multifractal systems generated from the multiplicative cascade and the MMAR methods. Without solving the associated eigenproblems, we calculated the landscape functions and accurately predicted the locations of the supported eigenmodes. Finally, based on the information obtained from the localization landscapes, we designed and fabricated multifractal photonic membranes in highly transparent SiN materials and directly imaged, using leaky-mode spectroscopy, their optical modes across the visible spectral range. The general predictions from the localization landscape of the Helmholtz operator were found to be in good qualitative agreement with the experimental data, establishing the localization landscape theory as a viable tool for the rapid exploration and benchmarking of scattering resonances in complex photonic structures with tailored multifractal disorder for nanophotonics and metamaterials applications. Finally, our results unveil the distinctive localization behavior of the scattering resonances supported by extended fractal and multifractal photonic membranes.